OpenStudy (anonymous):

What is the equation of the line, in standard form, that passes through (4, -3) and is parallel to the line whose equation is 4x + y - 2 = 0? 4x-y= 13?am i correct

OpenStudy (compassionate):

Show me your steps, Student.

OpenStudy (anonymous):

y=mx+b y=4x+0 y=4x

OpenStudy (anonymous):

Solve your equation for y to find the slope.

OpenStudy (anonymous):

idk how

OpenStudy (anonymous):

Everything is on the left side, you want everything except the y on the right side. So, use inverse operations to move the terms.

OpenStudy (anonymous):

\(4x + y - 2 = 0\) \(-4x ..... ..= -4x\)

OpenStudy (anonymous):

4x-2=y

OpenStudy (anonymous):

4 x 4=16-2=14

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

Subtract \(4x\) from each side.

OpenStudy (anonymous):

Then add 2 to each side.

OpenStudy (anonymous):

the equation was 4x + y - 2 = 0 i moved 4x and -2 to the left

OpenStudy (anonymous):

right i mean

OpenStudy (anonymous):

\(4x + y - 2 = 0\) Can be re-written as: \(y=-4x+2\) or \(4x-2 = -y\) which needs to be multiplied through by a -1 to get \(-4x+2=y\)

OpenStudy (anonymous):

Do you follow?

OpenStudy (anonymous):

4x-y=0?

OpenStudy (anonymous):

I think you're jumping ahead without understanding the process. I know you have a lot to do and limited time to do it, but please stay with me on this one. Ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The equation of the given line is \(y=−4x+2\) Do you understand how I got that?

OpenStudy (anonymous):

you multiplyed 4x-2=-y

OpenStudy (anonymous):

Good. Next, realize that a perpendicular line has a slope that is the negative inverse of the given line. So, what is the negative inverse of -4?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

The inverse means to flip the fraction. \[4 = \frac{4}{1}\]

OpenStudy (anonymous):

-1/4

OpenStudy (anonymous):

The inverse of -4 is \(\frac{1}{4}\)

OpenStudy (anonymous):

So, that is the slope of your new line. Use the point-slope formula to get the equation of the new line. \[y – y1 = m(x – x1)\]

OpenStudy (anonymous):

m= 1/4 Find \(x_1\) and \(y_1\) from the point given in the equation.

OpenStudy (anonymous):

\[y – y_1 = m(x – x_1)\]

OpenStudy (anonymous):

y=-1/4x+b?

OpenStudy (anonymous):

The slope is positive 1/4.

OpenStudy (anonymous):

Use the point-slope formula to get the equation: \(y – y_1 = m(x – x_1)\)

OpenStudy (anonymous):

y=1/4(4)+-3

OpenStudy (anonymous):

=13

OpenStudy (anonymous):

x and y will remain in the equation. You will substitute the numbers in your point for the \(x_1\) and \(y_1\).

OpenStudy (anonymous):

im not sure how to do that

OpenStudy (anonymous):

Using point (4,-3) and point-slope formula \(y – y_1 = m(x – x_1)\) You will get\( y+3 = \frac{1}{4}(x-4)\). Does that make sense?

OpenStudy (anonymous):

\(m= \frac{1}{4}\) \(y_1 = -3\) \(x_1 = 4\)

OpenStudy (anonymous):

i thnik i remember this can you show the equation includin x1 and y1 in a drawing?

OpenStudy (anonymous):

Not on this site. I don't know how. Let me see if I can find a better website....

OpenStudy (anonymous):

Try this: http://www.mathsisfun.com/data/function-grapher.php

OpenStudy (anonymous):

If the graph doesn't show up, just put the two equations in. y=-4x+2 y=1/4x+2

OpenStudy (anonymous):

im sorry im having a hard thime with this can we move on?

OpenStudy (anonymous):

Sure - as long as you promise me that you will continue to study after you complete the test so you will be ready for the next class.

OpenStudy (anonymous):

i will thank you